### v3

@@ 4.1 Dataset Construction
4.2 System Implementation
The OM4OV pipeline is implemented in Agent-OV, as an extension of Agent-OM [21]. Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [18] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [15], Alibaba Qwen [1], Google Gemma [7], and ChatGLM [25]. We use the OpenAI model gpt-4o-mini (2024-07-18) [17] in our experiments. For performance using different LLMs, we refer the reader to [21], where we find that API-accessed LLMs generally perform better than open-source LLMs. The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [8].
4.3 Experimental Results
Figure 4 shows the evaluation of Agent-OV and Agent-OM on the OV testbed. The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated. The results indicate that it is possible to unify the OM and OV tasks. Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 score provided. However, the original OM systems require necessary modifications to fit into the OV tasks. We can see that Agent-OM always has an unusually high overall performance across all alignments. This is because the traditional OM system only captures matched entities between two ontologies, where this refers to the remain and update entities in the OV task. While remain entities commonly have a large proportion between two different versions, the measure of remain entities dominates the overall evaluation. We believe this is a common pitfall and causes skewed performance measurement when using the OM system in OV tasks. Agent-OV overcomes the skewed measure in Agent-OM. When decomposing the measure into four sub-measures, we can observe a more precise matching performance across different OV tasks. In general, we observe the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. {bracketenumerate}
@@ 4.3 Experimental Results
Figure 5 shows the evaluation of using different similarity threshold methods in Agent-OV. We change the similarity threshold from 0.9 to 1 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similarly to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Furthermore, detecting update entities and detecting add and delete operations are negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities. Our explanations on these trends are as follows.
Figure 5: Evaluation of using different similarity threshold methods in Agent-OV.
@@
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$. For recall, there are direct relations between sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa. For precision, there are no direct relations between sub-measures. The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or either (?)?(?)( ? ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
$Rec\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}\quad Prec\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
$Rec\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}\quad Prec\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
@@ 5.3 Evaluation
Figure 7 shows the comparison of the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed.
Figure 8 shows the comparison of using different similarity thresholds in the CMT Ontology versioning. The results of the experiment also show that our cross-reference mechanism is less sensitive to the similarity threshold, which means it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds.
@@ 6 Discussion
False mappings in OV can result from a flawed ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
@@
False mappings in OV can create an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in the domain knowledge understanding.
@@
False mappings in OV can be an inappropriate setting of the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
@@ 8 Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks. Evaluations of our OAEI tested validate the feasibility and reliability of our system. Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, smart building ontologies Brick Schema [3], RealEstateCore [11], and ASHRAE Standard 223P [16].

### v4

@@ 4.2 Experiment Setup
The OM4OV pipeline is implemented in Agent-OV, as an extension of Agent-OM [19]. Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend Agent-OM with the OM4OV pipeline so that it can be used for OV tasks. Agent-OV supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [16] and Anthropic Claude [2], as well as open-source LLMs Meta Llama [14], Alibaba Qwen [1], Google Gemma [7], and ChatGLM [23]. For performance using different LLMs, we refer the reader to [19], where we find that API-accessed LLMs generally perform better than open-source LLMs. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV.
4.3 Results
Figure 4 evaluates Agent-OV and Agent-OM on the OV testbed. The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated. The results indicate that it is possible to unify the OM and OV tasks as developed in this paper. Both Agent-OM and Agent-OV can produce an alignment with the precision, recall, and F1 scores of Figure 4. However, the original OM systems require necessary modifications to tackle OV tasks. We can see that Agent-OM always has an unusually high overall performance across all alignments. This is because the traditional OM system only captures matched entities between two ontologies, where this refers to the remain and update entities in the OV task. While remain entities commonly comprise the major proportion of entities over two different versions, the measure of remain entities dominates the overall evaluation. We believe this is a common situation and causes misleading performance measurement when using the OM system in OV tasks. Agent-OV overcomes the skewed measure in Agent-OM. When decomposing the measure into four sub-measures, we can observe a more precise matching performance across different OV tasks. In general, we observe the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. {bracketenumerate}
@@ 4.3 Results
Figure 5 studies the effect of varying similarity thresholds in Agent-OV. We change the similarity threshold from 0.9 to 1 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similarly to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Furthermore, detecting update entities and detecting add and delete operations are negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities. Our explanations for these trends are as follows.
Figure 5: Evaluation varying similarity thresholds in Agent-OV. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
@@
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$. For recall, there are direct relations between sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa. For precision, there are no direct relations between sub-measures. The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or uncertain ($\sim$), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
$Rec\odot|\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|R|}\quad Prec\odot|\otimes\sim=\frac{|A\cap R|+|\Delta(A\cap R)\odot|\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
$Rec\oplus|\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|R|}\quad Prec\oplus|\ominus\sim=\frac{|A\cap R|-|\Delta(A\cap R)\oplus|\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
@@ 5.3 Evaluation
Figure 7 compares the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
Figure 7: Precision, Recall and F1 evaluation of the cross-reference (CR) mechanism on the OV testbed.
Figure 8 compares the effects of different similarity thresholds in the CMT Ontology versioning. The results of the experiment show that our cross-reference mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline, which means the cross-reference mechanism can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
@@ 6 Discussion
False mappings in OV can result from a flawed ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by inconsistent naming convention practice. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
@@
False mappings in OV can create an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in domain understanding or evolution of natural language in the domain.
@@
False mappings in OV can arise from an inappropriate setting of the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure each term is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
@@ 8 Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with minor modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks. Evaluation experiments on our OAEI testbed validate the feasibility and reliability of our system. Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods, that is, using URIs or additional versioning triples, and ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, smart building ontologies Brick Schema [3], RealEstateCore [10], and ASHRAE Standard 223P [15].